Flying against the jetstream, a jet travels 2920km in 4 hours. Flying with the jetstream, the same jet travels 7630km in 7 hours. What is the speed of the jet in still air, and what is the speed of the jetstream?

Found 2 solutions by Alan3354, checkley75:
Answer by Alan3354(69443)

Flying against the jetstream, a jet travels 2920km in 4 hours. Flying with the jetstream, the same jet travels 7630km in 7 hours. What is the speed of the jet in still air, and what is the speed of the jetstream?
----------------
2920 km/4 hrs = 730 kph ground speed
7630 km/7 hrs = 1090 kph ground speed
The airspeed of the plane is the average, (730+1090)/2
Airspeed = 910 kph
The wind speed is the difference between either groundspeed and the airspeed.
910 - 730 = 180 kph wind speed.
-------------
BTW, kilometers and kph are not used in aviation, we use nautical miles and knots. 1 knot is 1 nm/hour.
Flying with the wind is called downwind, or flying with a tailwind. Against it is called upwind, or into a headwind.

Answer by checkley75(3666)

D=RT
2920=R*4
R=2920/4
R=730 KMH AGAINST THE JETSTREAM.
7630=5*7
R=7630/7
R=1090 KMH QITH THE JETSTREAM.
(1090-730)/2=360/2=180 KMH IS THE SPEED OF THE JETSTREAM.
Proof:
730+180=910 speed of the jet in still air
1090-180=910 ditto.